Test the claim about the population mean, , at the given level of significance using the given sample statistics. Claim: u =% 7.2.38-T Question Help A company claims that the mean monthly residential electricity consumption in a certain region is moUse technology to help you test the claim about the population mean, y, at the given level of significance, a, using the giveTest the claim about the population mean, i, at the given level of significance using the given sample statistics. Claim: u #7.2.11 Find the P-value for the indicated hypothesis test with the given standardized test statistic, z. Decide whether to re
Answer #1

Solution : - Given that, n = 62 X = 28.8 6 3.02 The null and alternative hypothesis aze, Hoil = 30 Hand & 30 The test statista the null and alternative hypotheses are Ho u < 860 Hau > 860 Hese, u= 69 x = 880,0 = 120 the test statistic 2 z X-1 oldu TS& hypothesis aa. The null and alternative - Ho uc 1270 - Ha u z 1270 Here in= 300, x= 1285, o. 206.24 The test statistic z isGiven, test Statistic z = 1.83 - 2 = 0.03. Right tailed test - P-value = P(Z > 1.83] a 1-P(Z 2 1.83) 1 0.9664 - Using z tob..